The midpoint of the line segment ST has coordinates (3,6). If S is the point (4,1), find the coordinates of T.
I'm obviously stuck with this question though I did attempt to do it.
Here's my working out.
Let the coordinate of T be (x,y)
Then by the midpoint formula (x+3/2, x+6/2)=(4,1)
Midpoint formula for x.
x+3/2=4
x+3=8 (I've multiplied both sides by 4 though I don't quiet remember why now)
x=5
I haven't gone further than this onto the midpoint formula for y because I know that what I've done is completely wrong.
So I'd really appreciate it of someone did help me.
Thanks in advance =)

Answer by stanbon(75887)

The midpoint of the line segment ST has coordinates (3,6). If S is the point (4,1), find the coordinates of T.
I'm obviously stuck with this question though I did attempt to do it.
Here's my working out.
Let the coordinate of T be (x,y)
Then by the midpoint formula ((x+3)/2, (y+6)/2)=(4,1)
Midpoint formula for x.
(x+4)/2=3
x+4 = 6
x = 2
--------
(y+1)/2 = 6
y+1 = 12
y = 11
============
Note: Parentheses make a difference.
================
Cheers,
Stan H.